There are 20 rows of seats on a concert hall:25 seats are in the 1st row, 27 seats on the 2nd row,29 seats on the 3rd row,and so on. If the price per ticket is 2,300, how much will be the total sales for a one-night concert if all seats are taken
Answer by stanbon(75887)

There are 20 rows of seats on a concert hall:25 seats are in the 1st row, 27 seats on the 2nd row,29 seats on the 3rd row,and so on. If the price per ticket is 2,300, how much will be the total sales for a one-night concert if all seats are taken.
-----
Arithmetic Sequence with a(1) = 25 and d = 2
---
a(20) = 25 + 19*2 = 25 + 38 = 63
------
# of seats = (20/2)(25+63) = 10*88 = 880
----
Total revenue = 2.3*880 = $2024
------------
Cheers,
Stan H.
---------------